006 (part 1 of 2) 10.0 points Assume: The system shown below starts from rest and there are no frictional forces opposing the motion of the system. The dolly and its wheels have negligible mass. The acceleration of gravity is 9.8 m/s 23.8 N 281 N of If a person pulls on the rope with a constant force what is the acceleration of the system? Answer in units of m/s? 007 (part 2 of 2) 10.0 points How far will it move in 3.16 s? Answer in units of m. era-
Answer #1

Solution,

Given,

Weight, W = 281 N

Force F = 23.8 N

Mass, m = W/g = 281/9.8 = 28.67 kg

Acceleration, a = F/m = 23.8/28.67 = 0.83 m/s^2  (Ans)

Distance, d = 0.5 at^2 = 0.5 x 0.83 x 3.16^2 = 4.14 m  (Ans)

Comment in case any doubt please rate my answer ....